OpenStudy (anonymous):

rewrite the equation 1 + cos x -2 sin^2 x = 0 in terms of cos x.

OpenStudy (mathstudent55):

\(\sin^2 x + \cos^2 x = 1\) Solve for \( \sin^2 x\) and substitute.

OpenStudy (rock_mit182):

remember that :\[\sin ^{2} + \cos ^{2}=1\]

OpenStudy (anonymous):

i dont know how to do that and want an actual answer

myininaya (myininaya):

Can you solve the equation math gave you for sin^2 This equation sin^2(x)+cos^2(x)=1? Can you get sin^2(x) by itself here?

myininaya (myininaya):

 how do we undo that addition of cos^2(x) to obtain just sin^2(x) on one side by itself.